# Batched Nonparametric Contextual Bandits

### v2

@@ B.2 Regret decomposition
$R_{T}(\pi)$ $=r_{T}^{\textrm{born}}(\mathcal{X})$
$=\underbrace{r_{T}^{\textrm{live}}(\mathcal{X})}_{=0}+\sum_{C^{\prime}\in\mathsf{child}(\mathcal{X})}r_{T}^{\textrm{born}}(C^{\prime})$
$=\sum_{1\leq i<M}\left(\underbrace{\sum_{C\in\mathcal{B}_{i}}r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})}_{\eqqcolon U_{i}}+\underbrace{\sum_{C\in\mathcal{B}_{i}}r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})}_{\eqqcolon V_{i}}\right)$
$\quad+\sum_{C\in\mathcal{B}_{M}}r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}),$
@@ B.4 Putting things together
$\mathbb{E}[R_{T}(\pi)]$ $\leq c\left(t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}\right),$
@@ C.2 Three failure cases for $g$
$\mathbb{E}[R_{T}(\hat{\pi})]$ $\overset{\mathrm{(i)}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@
$\mathbb{E}[R_{T}(\hat{\pi})]$ $\geq\sum_{t=t_{1}+1}^{t_{2}}\sum_{l=H/4}^{3H/4}\mathbb{E}\left[\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\mathbf{1}\{X_{t}\in B_{l}\}\right]$
@@
$\mathbb{E}[R_{T}(\hat{\pi})]$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@
$\mathbb{E}[R_{T}(\hat{\pi})]$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{2}+1}^{T}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@ Appendix D Proof of Theorem 5
When $\beta=d=1$, we denote $\gamma(\alpha^{\star})=(\alpha^{\star}+1)/3$. Fix some small constant $\epsilon>0$ to be specified later. We deal with $M=2$ and $M=3$ separately. In both cases, we use the fact that the algorithm needs to provide its first batch size $t_{1}$ prior to the game, and design $\alpha^{\star}$ such that any choice of $t_{1}$ would fail.
@@ D.1 When $M=2$
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq t_{1}=T^{\frac{3}{5}+\kappa_{1}},$
where $\kappa_{1}=\epsilon$.
@@
Take $\alpha^{\star}=o(1)$. By relation (9), we have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}=T^{1-(\frac{3}{4}-\epsilon)\frac{\alpha^{\star}+1}{3}}=T^{\frac{3}{4}+\kappa_{1}},$
where $\kappa_{1}=(\alpha^{\star}+1)\epsilon/3-\alpha^{\star}/4>0$.
D.2 When $M=3$
Under $M=3$, the theoretical optimal rate is $T^{(1-\gamma(\alpha^{\star}))/(1-\gamma(\alpha^{\star})^{3})}=T^{9/((\alpha^{\star})^{2}+5\alpha^{\star}+13)}$.
Case of $t_{1}=T^{9/19+\epsilon}$.
@@ D.2 When $M=3$
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim t_{1}=T^{\frac{9}{19}+\kappa_{1}},$
where $\kappa_{1}=\epsilon$.
Case of $t_{1}=T^{9/13-\epsilon}$.
Take $\alpha^{\star}=o(1)$. Denote $T_{2}=T^{1/(\gamma(\alpha^{\star})+1)}t_{1}^{\gamma(\alpha^{\star})/(\gamma(\alpha^{\star})+1)}$. Define the events $E_{2}=\{T_{2}<t_{2}\}$ and $E_{3}=\{t_{2}\leq T_{2}\}$. Recall $Q_{i}(\cdot)=\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)\mathbb{P}_{\pi,i,\omega}(\cdot)$ from (13) given $z_{i}>0$ and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. Take $z_{2}=\lceil(36t_{1}2^{2})^{1/(2\beta+d)}\rceil$ and $z_{3}=\lceil(36T_{2}2^{2})^{1/(2\beta+d)}\rceil$. Since $E_{2}$ can be determined by observations up to $t_{1}$, we have
$|Q_{2}(E_{2})-Q_{3}(E_{2})|=|Q_{2}^{t_{1}}(E_{2})-Q_{3}^{t_{1}}(E_{2})|\leq\mathrm{TV}(Q_{2}^{t_{1}},Q_{3}^{t_{1}})\overset{\mathrm{(i)}}{\leq}\frac{1}{2}\sqrt{t_{1}z_{2}^{-(2\beta+d)}}\overset{\mathrm{(ii)}}{\leq}\frac{1}{4},$ (21)
where step (i) applies Lemma 2, and step (ii) is due to the definition of $z_{2}$. Consequently,
$Q_{2}(E_{2})+Q_{3}(E_{3})$ $=Q_{2}(E_{2})-Q_{3}(E_{2})+Q_{3}(E_{2})+Q_{3}(E_{3})$
$\geq-\frac{1}{4}+Q_{3}(E_{2})+Q_{3}(E_{3})=\frac{3}{4},$
where the second step uses inequality (21) and the last equality is because $E_{2}$ and $E_{3}$ form a whole partition of the probability space. Then we would have at least one of $Q_{2}(E_{2})\geq 1/4$ or $Q_{3}(E_{3})\geq 1/4$. If $Q_{2}(E_{2})\geq 1/4$, by Lemma 4 we obtain
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\geq\sup_{f\in\mathcal{C}_{z_{2}}}R_{T_{2}}(\pi;f)\apprge T_{2}z_{2}^{-\beta(1+\alpha^{\star})}\asymp T^{\frac{9}{13}+\kappa_{1}},$
for some $\kappa_{1}>0$. If $Q_{3}(E_{3})\geq 1/4$, we similarly have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\geq\sup_{f\in\mathcal{C}_{z_{3}}}R_{T_{3}}(\pi;f)\apprge Tz_{3}^{-\beta(1+\alpha^{\star})}\asymp T^{\frac{9}{13}+\kappa_{1}},$
for some $\kappa_{1}>0$.

### v3

@@ B.2 Regret decomposition
$R_{T}(\pi)$ $=\mathbb{E}[r_{T}^{\textrm{born}}(\mathcal{X})]$
$=\underbrace{\mathbb{E}[r_{T}^{\textrm{live}}(\mathcal{X})]}_{=0}+\sum_{C^{\prime}\in\mathsf{child}(\mathcal{X})}\mathbb{E}[r_{T}^{\textrm{born}}(C^{\prime})]$
$=\sum_{1\leq i<M}\left(\underbrace{\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})]}_{\eqqcolon U_{i}}+\underbrace{\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})]}_{\eqqcolon V_{i}}\right)$
$\quad+\sum_{C\in\mathcal{B}_{M}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C})],$
@@ B.4 Putting things together
$R_{T}(\pi)$ $\leq c\left(t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}\right),$
@@ C.2 Three failure cases for $g$
$R_{T}(\hat{\pi})$ $\overset{\mathrm{(i)}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@
$R_{T}(\hat{\pi})$ $\geq\sum_{t=t_{1}+1}^{t_{2}}\sum_{l=H/4}^{3H/4}\mathbb{E}\left[\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\mathbf{1}\{X_{t}\in B_{l}\}\right]$
@@
$R_{T}(\hat{\pi})$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@
$R_{T}(\hat{\pi})$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{2}+1}^{T}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@ Appendix D Proof of Theorem 5
Recall $\gamma(\alpha^{\star})=\beta(\alpha^{\star}+1)/(2\beta+d)=(\alpha^{\star}+1)/3=\gamma^{\star}$. Fix some small constant $\epsilon>0$. We first deal with $M=2$ and then extend the analysis to general $M$. Throughout the proof, we use the fact that the algorithm needs to provide its first batch size $t_{1}$ prior to the game, and design $\alpha^{\star}$ such that any choice of $t_{1}$ would fail.
@@ D.1 When $M=2$
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi)\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq t_{1}=T^{\frac{3}{5}+\epsilon}.$
@@
Take $\alpha^{\star}>0$ such that $T^{3/(\alpha^{\star}+4)}>t_{1}$. By relation (9), we have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi)\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}=T^{\frac{3}{\alpha^{\star}+4}+(\alpha^{\star}+1)(\frac{\epsilon}{3}+\frac{1}{\alpha^{\star}+4}-\frac{1}{4})},$
where $\frac{\epsilon}{3}+\frac{1}{\alpha^{\star}+4}-\frac{1}{4}>0$ due to the choice of $\alpha^{\star}$.
D.2 General $M$
Since $(1-\gamma(\alpha))/(1-\gamma^{M}(\alpha))$ is decreasing as $\alpha$ increases, one has
$\frac{1-\gamma(1)}{1-\gamma^{M}(1)}<\frac{1-\gamma(0)}{1-\gamma^{M}(0)}.$
We can divide the proof into the below two cases.
Case of $t_{1}=T^{(1-\gamma(1))/(1-\gamma^{M}(1))+\epsilon}$.
@@ D.2 General $M$
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi)\gtrsim t_{1}=T^{\frac{1-\gamma^{\star}}{1-(\gamma^{\star})^{M}}+\epsilon}.$
Case of $t_{1}=T^{(1-\gamma(0))/(1-\gamma^{M}(0))-\epsilon}$.
Take $\alpha^{\star}>0$ such that $T_{1}=T^{(1-\gamma^{\star})/(1-(\gamma^{\star})^{M})}>t_{1}$. Define
$T_{2}=T^{\frac{1-\gamma^{\star}}{1-(\gamma^{\star})^{M-1}}}\cdot t_{1}^{\frac{\gamma^{\star}-(\gamma^{\star})^{M-1}}{1-(\gamma^{\star})^{M-1}}},$
and
$T_{i}=t_{1}^{\frac{\gamma^{\star}-(\gamma^{\star})^{i-1}}{\gamma^{\star}-1}}\cdot T_{2}^{\frac{(\gamma^{\star})^{i-1}-1}{\gamma^{\star}-1}}$
when $3\leq i\leq M$. It can be verified that $T_{M}=T$. Define the events
$A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}.$
for $2\leq i\leq M$. Moreover, take $z_{2}=\lceil(36t_{1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ and $z_{i}=\lceil(36T_{i-1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ for $3\leq i\leq M$. Recall $Q_{i}(\cdot)=\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)\mathbb{P}_{\pi,i,\omega}(\cdot)$ from (13) given $z_{i}>0$ and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. We can apply Lemma 3 with $k=2$ to obtain that there exists $i^{\star}\in\{2,...,M\}$ such that $Q_{i^{\star}}(A_{i^{\star}})\geq 1/(2(M-1))$. Consequently, by Lemma 4 we reach
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi;f)$ $\geq\sup_{f\in\mathcal{C}_{z_{i^{\star}}}}R_{T_{i^{\star}}}(\pi;f)$
$\apprge T_{i^{\star}}z_{i^{\star}}^{-\beta(1+\alpha^{\star})}(M-1)^{-\frac{1+\alpha^{\star}}{\alpha^{\star}}}$
$\asymp T^{\frac{1-\gamma^{\star}}{1-(\gamma^{\star})^{M}}\;+\;c_{2}(\gamma^{\star})^{M-1}\cdot\frac{1-\gamma^{\star}}{1-(\gamma^{\star})^{M-1}}}\cdot(\frac{1}{M})^{c_{3}},$
for some $c_{2},c_{3}>0$.